If Indiana stayed the same size in area, but had the same population density as Alaska,...

  1. If Indiana stayed the same size in area, but had the same population density as Alaska, how many people would we expect to live in Indiana? How does this compare to the actual population of Indiana?
Answer #1

Area of Indiana is 94,322 sq.km.

population density of alaska is 0.49 per sq.km

now the population if the density is same as alaska is = 94,322 * 0.46 = 46217.78 ~ 46218

now actual population of Indiana is 6691878

6691878/46218 = 144.789 ~ 144.8

hence it would be less than 144.8 times less than the actual population of Indiana